# Multistage Collaborative Knowledge Distillation from a Large Language Model for Semi-Supervised Sequence Generation

### v3

@@ Abstract
We study semi-supervised sequence generation tasks, where the few labeled examples are too scarce to finetune a model, and meanwhile, few-shot prompted large language models (LLMs) exhibit room for improvement. In this paper, we present the discovery that a student model distilled from a few-shot prompted LLM can commonly generalize better than its teacher to unseen examples on such tasks. We find that the student is able to learn a general pattern from the high-quality pseudolabels produced by the teacher during knowledge distillation (KD), and favorably not a general pattern from the low-quality pseudolables. Leveraging this discovery, we propose a new method, Multistage Collaborative Knowledge Distillation from an LLM (MCKD), for these tasks. MCKD first few-shot prompts an LLM to produce pseudolabels for unlabeled data. Then at each stage of an iterative KD process, a new pair of students is trained on disjoint partitions of the pseudolabeled data, and produces new and improved pseudolabels for their unseen partitions. We conduct extensive experiments on four syntactic and semantic parsing datasets and show the effectiveness of MCKD for low-resource semi-supervised sequence generation. On CRAFT biomedical parsing, for example, 3-stage MCKD with 50 labeled examples outperforms an LLM teacher and vanilla KD by 7.5% and 3.7% parsing F1, respectively, and matches the performance of supervised finetuning with 500 labeled examples.
@@ 1 Introduction
Leveraging the discovery, we propose a novel method, Multistage Collaborative KD from an LLM (MCKD), for semi-supervised sequence generation. MCKD first collects pseudolabels for a large amount of unlabeled data from a few-shot prompted LLM, bootstraping a multistage KD process. At each KD stage, a new pair of students is trained on distinct partitions of the pseudolabeled data and asked to produce pseudolabels for the data that they have not been trained on. This process improves upon vanilla KD through two core mechanisms: (1) collaborative cross-partition labeling at each KD stage, which splits all pseudolabeled data into two mutually exclusive partitions to train a collaborative pair of students and leverages their strong generalization capabilities to relabel the data, and (2) multistage KD, whereby student models continue to generalize better and produce higher quality pseudolabels than their teachers over multiple KD stages.
@@ 3.1 Data and evaluation
We compute token-level F1-score for each example and average over examples to measure model performance. For constituency parsing, we follow common practice to use bracketing but discard constituent tags in training and evaluation, and remove punctuation before computing parsing F1.333Decisions associated with punctuation are often arbitrary in ground truth parse trees and not informative of model performance. We minimally post-process model outputs, balancing brackets and fixing accidental word segmentation. [^4] We do not constrain model generation. Early during training, T5 fails to consistently output well-formed parse trees. This is less of an issue after the first few epochs. For semantic parsing, we use the evaluation function from Goo et al. (2018). [^5] https://github.com/MiuLab/SlotGated-SLU/blob/master/utils.py
@@ 3.2 Model, learning, and prediction
We use GPT-3.5 Turbo [^6] https://platform.openai.com/docs/models, a model supported by the ChatGPT API, to collect initial pseudolabels for $\mathcal{D}^{\text{unlabeled}}$. We access the model through the OpenAI API. We randomly and uniformly sample 30 labeled parsing examples from $\mathcal{D}^{\text{labeled}}$ to include in the prompt, since 50 or 250 examples will exceed the allowed maximum context length. Details of prompt design can be found in Appendix C.
@@
Our student models are finetuned from T5-Base (220M parameters; Raffel et al., 2020). Our assumed low-resource semi-supervised setting implies that the amount of available labeled data is small and a reliably large validation set does not exist. Inspired by classic unsupervised algorithms such as K-Means, we train each student until a convergence criteria is met — we train the student until the training accuracy on teacher pseudolabels changes by less than 0.1% for three consecutive epochs. Generation is performed via greedy decoding. The other hyperparameters are learning rate $3\times 10^{-4}$, batch size 32, and $21$ maximum training epochs. [^7] Oracle experiments on PTB show that the validation performance doesn’t change significantly given a wide range of these hyperparameters, and thus we use a default set in all experiments. In scenarios where a labeled validation set is available, further tuning may prove even more beneficial due to optimization behavior such as double descent Opper et al. (1989); LeCun et al. (1991); Maddox et al. (2020). Results are from single runs unless otherwise stated.
@@ 4 Main Results
As shown in Table 1, GPT-3.5 Turbo’s few-shot performance is reasonably good but has room for improvement. Due to the maximum input length for GPT-3.5 Turbo, we always sample 30 in-context demonstrations. In most cases, sampling them from a set of 50 of 250 labeled data $\mathcal{D}^{\text{labeled}}$ does not make a big difference.
@@ 5.2 The student corrects the teacher’s errors
5.3 Large distillation sets are useful
A stage-2 student $s_{2}$ is then trained similarly on $\mathcal{D}_{B}$ with pseudolabels from $s_{1}$ and evaluated on the test set. We examine the effect of the sizes of distillation sets $\mathcal{D}_{A}$ and $\mathcal{D}_{B}$ on the final test performance of $s_{2}$ in Figure 7. When the distillation set $\mathcal{D}_{A}$ for $s_{1}$ is too small, increasing the size of $\mathcal{D}_{B}$ by generating more pseudolabels with $s_{1}$ does not substantially improve performance for $s_{2}$, likely due to the pseudolabels being too noisy. Learning from an increasing amount of such noisy labels may even worsen the test set performance of $s_{2}$; for instance, with a $\mathcal{D}_{A}$ of 2k instances, enlarging $\mathcal{D}_{B}$ from 4k to 8k instances reduces the test F1 of $s_{2}$ from 71.3 to 69.0. On the other hand, when $\mathcal{D}_{B}$ is too small, making $\mathcal{D}_{A}$ larger alone does not improve the test set performance of $s_{2}$ either, despite higher-quality pseudolabels in $\mathcal{D}_{B}$ from a better-trained $s_{1}$. Overall, the above results indicate that adequately large distillation sets are needed at every stage of MCKD to optimize the performance of the final student. In the original, more streamlined implementation in Section 2, we employ a pair of students at each stage to pseudolabel examples, so that the final student can take advantage of the most data available.
@@ 7 Conclusion
Semi-supervised sequence generation in specialized low-resource settings can be challenging for NLP systems. This paper presents a novel solution to this problem that is able to leverage an LLM for training much smaller yet higher-accuracy models, despite the LLM itself performing only at a mediocre level. We first present the discovery that student models distilled from a few-shot prompted LLM can often correct the teacher’s errors, thus showing superior generalization, on such tasks. Exploiting this phenomenon along with strategies such as data partitioning and cross-partition labeling over multiple stages, we then propose MCKD, a novel approach that utilizes knowledge distillation (KD) to train high-accuracy students. MCKD essentially leverages progressively better models to train future generations with even better performance. Extensive empirical evaluation verifies the effectiveness and label efficiency of MCKD.
@@ Appendix A Dataset Details
In Table 3, we show examples of the contituency and semantic parsing tasks that we work with. We use the training and testing sets described in 3.1. We assume a low-resource scenario where a labeled validation set is unavailable; we find default parameters that are not tuned on dataset-specific validation sets are effective across datasets. For PTB we adopt the standard data splits.
@@
The standard split for the test set of PTB contains 2416 examples. For training, we randomly sample 20,000 examples as unlabeled training data. CRAFT contains 21,121 examples, but does not have standard splits; we randomly sample 50% examples as the training set and 40% examples as the test set. The remaining could be used for validation, although we assume a low-resource scenario where a labeled validation set is unavailable. For CRAFT, the test set contains 8448 examples, while the unlabeled training set contains 10560 examples.
@@
Appendix B Implementation Details
Hardware and software.
Each experiment takes less than 6 hours on a single RTX 6000 GPU. Code is based on huggingface transformers Wolf et al. (2020) and will be open-sourced.
Appendix C Prompting LLM for Pseudolabel Generation
@@
Appendix D Effects of Training Fidelity in Distillation Sets of Different Sizes
As discussed in Section 5.1, we find high training fidelity, can generally lead to a better student’s performance on held-out data, despite the unreliability of teacher labels. This section investigates whether the observation still holds with different scales of the distillation set, especially when the size is substantially smaller. Results are shown in Figure 8. Similar patterns across different scales of distillation sets can be observed. Even if the distillation set only contains 250 teacher labels, the student can still generally have better performance on held-out data rather when the student memorizes more teacher labels.
Table 3: Examples from our used datasets.

### v4

@@ Abstract
We study semi-supervised sequence generation tasks, where the few labeled examples are too scarce to finetune a model, and meanwhile, few-shot prompted large language models (LLMs) exhibit room for improvement. In this paper, we present the discovery that a student model distilled from a few-shot prompted LLM can commonly generalize better than its teacher to unseen examples on such tasks. We find that the student is able to learn a general pattern from the high-quality pseudolabels produced by the teacher during knowledge distillation (KD), and favorably not a general pattern from the low-quality pseudolabels. Leveraging this discovery, we propose a new method, Multistage Collaborative Knowledge Distillation from an LLM (MCKD), for these tasks. MCKD first few-shot prompts an LLM to produce pseudolabels for unlabeled data. Then at each stage of an iterative KD process, a new pair of students is trained on disjoint partitions of the pseudolabeled data, and produces new and improved pseudolabels for their unseen partitions. We conduct extensive experiments on four syntactic and semantic parsing datasets and show the effectiveness of MCKD for low-resource semi-supervised sequence generation. On CRAFT biomedical parsing, for example, 3-stage MCKD with 50 labeled examples outperforms an LLM teacher and vanilla KD by 7.5% and 3.7% parsing F1, respectively, and matches the performance of supervised finetuning with 500 labeled examples.  [^1] Our code is made available github.com/andotalao24/Multistage-Collaborative-Knowledge-Distillation.
@@ 1 Introduction
Leveraging the discovery, we propose a novel method, Multistage Collaborative KD from an LLM (MCKD), for semi-supervised sequence generation. MCKD first collects pseudolabels for a large amount of unlabeled data from a few-shot prompted LLM, bootstraping a multistage KD process. At each KD stage, a new pair of students is trained on distinct partitions of the pseudolabeled data and asked to produce pseudolabels for the data that they have not been trained on. This process improves upon vanilla KD through two core mechanisms: (1) cross-partition labeling at each KD stage, which splits all pseudolabeled data into two mutually exclusive partitions to train a collaborative pair of students and leverages their strong generalization capabilities to relabel the data, and (2) multistage KD, whereby student models continue to generalize better and produce higher quality pseudolabels than their teachers over multiple KD stages.
@@ 3.1 Data and evaluation
We compute token-level F1-score for each example and average over examples to measure model performance. For constituency parsing, we follow common practice to use bracketing but discard constituent tags in training and evaluation, and remove punctuation before computing parsing F1.444Decisions associated with punctuation are often arbitrary in ground truth parse trees and not informative of model performance. We minimally post-process model outputs, balancing brackets and fixing accidental word segmentation. [^5] We do not constrain model generation. Early during training, T5 fails to consistently output well-formed parse trees. This is less of an issue after the first few epochs. For semantic parsing, we use the evaluation function from Goo et al. (2018). [^6] https://github.com/MiuLab/SlotGated-SLU/blob/master/utils.py
@@ 3.2 Model, learning, and prediction
We use GPT-3.5 Turbo [^7] https://platform.openai.com/docs/models, a model supported by the ChatGPT API, to collect initial pseudolabels for $\mathcal{D}^{\text{unlabeled}}$. We access the model through the OpenAI API. We randomly and uniformly sample 30 labeled parsing examples from $\mathcal{D}^{\text{labeled}}$ to include in the prompt, since 50 or 250 examples will exceed the allowed maximum context length. Details of prompt design can be found in Appendix B.
@@
We finetune T5-Base (220M parameters; Raffel et al., 2020) as the student model using the Huggingface transformers library Wolf et al. (2020). Our assumed low-resource semi-supervised setting implies that the amount of available labeled data is small and a reliably large validation set does not exist. Inspired by classic unsupervised algorithms such as K-Means, we train each student until a convergence criteria is met — we train the student until the training accuracy on teacher pseudolabels changes by less than 0.1% for three consecutive epochs. The other hyperparameters are learning rate $3\times 10^{-4}$, batch size 32, and $21$ maximum training epochs. [^8] Oracle experiments on PTB show that the validation performance doesn’t change significantly given a wide range of these hyperparameters, and thus we use a default set in all experiments. In scenarios where a labeled validation set is available, further tuning may prove even more beneficial due to optimization behavior such as double descent Opper et al. (1989); LeCun et al. (1991); Maddox et al. (2020). Generation is performed via greedy decoding. Results are from single runs that take less than 6 hours on a single RTX 6000 GPU unless otherwise stated.
@@ 4 Main Results
As shown in Table 1, GPT-3.5 Turbo’s few-shot performance is reasonably good but has room for improvement. Due to the maximum input length for GPT-3.5 Turbo, we always sample 30 in-context demonstrations. In most cases, sampling them from a set of 50 or 250 labeled data $\mathcal{D}^{\text{labeled}}$ does not make a big difference.
@@ 5.2 The student corrects the teacher’s errors
5.3 Large distillation sets are useful
A stage-2 student $s_{2}$ is then trained on $\mathcal{D}_{B}$ with pseudolabels from $s_{1}$ and evaluated on the test set. We examine the effect of the sizes of distillation sets $\mathcal{D}_{A}$ and $\mathcal{D}_{B}$ on the final test performance of $s_{2}$ in Figure 7. When the distillation set $\mathcal{D}_{A}$ for $s_{1}$ is too small, increasing the size of $\mathcal{D}_{B}$ by generating more pseudolabels with $s_{1}$ does not substantially improve performance for $s_{2}$, likely due to the pseudolabels being too noisy. Learning from an increasing amount of such noisy labels may even worsen the test set performance of $s_{2}$; for instance, with a $\mathcal{D}_{A}$ of 2k instances, enlarging $\mathcal{D}_{B}$ from 4k to 8k instances reduces the test F1 of $s_{2}$ from 71.3 to 69.0. On the other hand, when $\mathcal{D}_{B}$ is too small, making $\mathcal{D}_{A}$ larger alone does not improve the test set performance of $s_{2}$ either, despite higher-quality pseudolabels in $\mathcal{D}_{B}$ from a better-trained $s_{1}$. Overall, the above results indicate that adequately large distillation sets are needed at every stage of MCKD to optimize the performance of the final student. In the original, more streamlined implementation in Section 2, we employ a pair of students at each stage to pseudolabel examples, so that the final student can take advantage of the most data available.
@@ 7 Conclusion
Semi-supervised sequence generation in specialized low-resource settings can be challenging for NLP systems. We present the discovery that student models distilled from a few-shot prompted LLM can often correct the teacher’s errors, thus showing superior generalization, on such tasks. Exploiting this phenomenon along with strategies such as data partitioning and cross-partition labeling over multiple stages, we then propose MCKD, a novel solution that utilizes knowledge distillation (KD) from an LLM, whose performance can be only at a mediocre level, to train much smaller yet high-accuracy students. MCKD essentially leverages progressively better models to train future generations with even better performance. Extensive empirical evaluation verifies the effectiveness and label efficiency of MCKD.
@@ Appendix A Dataset Details
In Table 2, we show examples of the contituency and semantic parsing tasks that we work with. We use the training and testing sets described in Section 3.1. We assume a low-resource scenario where a labeled validation set is unavailable; we find default parameters that are not tuned on dataset-specific validation sets are effective across datasets. For PTB we adopt the standard data splits.
@@
The standard split for the test set of PTB contains 2,416 examples. For training, we randomly sample 20,000 examples as unlabeled training data. CRAFT contains 21,121 examples, but does not have standard splits; we randomly sample 50% examples as the training set and 40% examples as the test set. The remaining could be used for validation, although we assume a low-resource scenario where a labeled validation set is unavailable. For CRAFT, the test set contains 8,448 examples, while the unlabeled training set contains 10,560 examples.
@@
Table 2: Examples from our used datasets.
Appendix B Prompting an LLM Teacher for Pseudolabel Generation
@@
Input:{input string of exemplar $i$}
Output:{output result of exemplar $i$}
Input:{input string of test data}
Output:
Table 3: The prompting format for LLM.
Appendix C Statistical Significance of Results
We present the statistical significance of the results in Table 1. On each dataset, for the setting that uses 50 labeled examples, we perform a paired t-test to compare (1) the distribution of per-example F1 scores of predictions by the best KD baseline and (2) that of predictions by the 2-stage MCKD. Under the null hypothesis, the distribution of the per-example F1 scores achieved by the strongest baseline model is the same as the per-example F1 scores achieved by the 2-stage MCKD. In Table 4, we observe that the p-values are all far below the commonly used 0.05 threshold. Therefore, we reject the null hypotheses and conclude that the 2-stage MCKD is significantly better than the strongest baselines in the mentioned dataset settings.
Dataset PTB CRAFT ATIS Snips
Best baseline KD+SD (top 25%) KD+SD (top 75%) KD+SD KD+SD
p-value 5.7e-10 5.3e-22 9.9e-5 1.2e-3
Table 4: P-values of paired t-tests between the per-example F1 distribution of our 2-stage MCKD and that of the best baselines. We observe that the MCKD is significantly better than the other approaches across different datasets.
Appendix D Relation Between Training Fidelity and Generalization Ability
As discussed in Section 5.1, we find high training fidelity of a student model is generally correlated with better performance on held-out data, despite the unreliability of teacher labels. This section investigates whether the observation still holds with different scales of the distillation set, especially when the size is substantially smaller. Results are shown in Figure 8. Similar patterns across different scales of distillation sets can be observed. Even if the distillation set only contains 250 teacher pseudolabels, the observation holds true.